OpenStudy (anonymous):

find the general solution of the differential equation dy/dx=x^2+2/3y^2

terenzreignz (terenzreignz):

like this \[\huge \frac{dy}{dx}=\frac{x^2+2}{3y^2}\]?

OpenStudy (anonymous):

It could also be the case that you meant \[\frac{dy}{dx}=x^2+\frac{2}{3}y^2\] or \[\frac{dy}{dx}=x^2+\frac{2}{3y^2}\] Both of which (I believe) are Bernoulli equations and can be solved in a similar fashion.

terenzreignz (terenzreignz):

@SithsAndGiggles I sure do hate ambiguities :)